OpenStudy (anonymous):

find the x and y intercepts for the graph equation: 6x+3y=-5 PLEASE HELP ME

sam (.sam.):

\[y=-2 x-\frac{5}{3}\]

OpenStudy (anonymous):

when x =0( y intercept): 0 + 3y =-5 y=-5/3 when y=0 (x intercept): 6x +0 =-5 x = -5/6

sam (.sam.):